Speculator considers purchasing several PUT options for a delivery of 100 kg of moon powder each with a strike price of 800 CHF per kg and premium of 400 CHF per contract. How many options at max can he enter if he owns 10,000 CHF? What would be his overall result, if the market price at delivery date is 860 CHF per kg?
Answer #1

Put Option is a derivative which provides its holder a right not the obligation to sell underlying assets at specified price on expiration date of option.

In above case -

Each put option contract - 100 kg of moon powder

Strike price(exercise price) = 800 CHF per kg

Put option premium = 400 CHF per contract.

No. of Put option contract can be purchased if he own 10,000 CHF

No of Put option contract = 10,000/400

= 25 put option contracts

Put option pay-off = MAX(Strike Price - Market Price , 0)

Now, if market Price at delivery date is 860 CHF

Put Option Pay-off = 100*25*{MAX(800-860,0)

= 0

Investor Profit/Loss = Put option pay-off - Put premium

= 0 - 400*25

= - 10,000 CHF

Speculator result on delivery date would be loss of 10,000 CHF.